A water drop of radius 10^-5cm is given a charge 1.6× 10^-19c.If it floats in a electric field,what is the intensity of the electric field?
1
Expert's answer
2020-06-08T10:25:30-0400

some important formula used here

mass = volume * density

density "=\frac43\pi r^3"

Force by electric field = Force of gravity

"1.6\times10^{-19}\times E=\frac{4\pi\times(10^{-7})^3}{3}\times1000\times10"

"E=2.61\times10^{2}" volt/ m